# SAP-X2C: Optimally-Simple Two-Component Relativistic Hamiltonian With Size-Intensive Picture Change

### v1

@@ Abstract
We present a simple relativistic exact 2-component (X2C) Hamiltonian that models two-electron picture-change effects using Lehtola’s superposition of atomic potentials (SAP) [S. Lehtola, J. Chem. Theory Comput. 15, 1593–1604 (2019)]. The SAP-X2C approach keeps the low-cost and technical simplicity of the popular 1-electron X2C (1eX2C) predecessor, but is significantly more accurate and has a well-defined thermodynamic limit, making it applicable to extended systems (such as periodic crystals). The assessment of the SAP-X2C-based Hartree-Fock total and spinor energies, spin-orbit splittings, equilibrium bond distances, and harmonic vibrational frequencies suggests that SAP-X2C is similar to the more complex atomic mean-field (AMF) X2C counterparts in the ability to approximate the 4-component Dirac-Hartree-Fock reference.
@@ 1 Introduction
Relativistic quantum chemistry1, 2, 3 is an approximation of quantum electrodynamics4 that combines Dirac 1-particle kinematics with the classical description of electromagnetic interactions in a many-particle system. By accounting for the dominant relativistic effects that are relevant for the chemistry and low-energy physics of atomistic matter and preserving the structure of nonrelativistic quantum many-body theory, accurate practical applications have become possible1, 5, 6, 7, 8, 9, 2, 10, 11, 12, 13, 14, 15. Despite the close resemblance to the nonrelativistic quantum chemistry, the appearance of positron-like degrees of freedom in the theory, such as bispinors (instead of spinors) and negative energy states, complicates both formal and practical aspects of relativistic quantum chemistry. Thus, from the earliest days, there was keen interest in constructing electron-only (2-component, 2C) heuristics that accurately approximate the reference 4-component (4C) treatment. So-called quasi-relativistic 2C methods16, 17, 18, 10, 19, 20, 21, 22, 23, 24 include: the Foldy-Wouthuysen (FW)25 transformation, the Douglas-Kroll-Hess (DKH)26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36 methods, the zeroth-order regular approximation (ZORA)37, 38, 38, 39, Barysz-Sadlej-Snijders (BSS)40, 19, 41 method, and exact-two-component (X2C)42 methods, which are generalized forms of the one-step normalized elimination of small component (NESC) approach prescribed by Dyall43, 44, 45, 46, 47. Our work adds to the X2C family of methods48, 49, 50, 51, 52, 53, whose common trait is that the Fock-space Hamiltonian $H$ and other property operators expressed in a finite basis are rotated by a unitary matrix $U$ that is designed to block diagonalize the 1-particle Hamiltonian $h$ or its effective Fock-like equivalent (see Appendix A for notation). Subsequently, the rotated Hamiltonian is projected onto the upper (electron-like) components of the bispinor:
@@
The myriad X2C variants differ in the details of these two design factors; see reviews by Liu18 and Wang et. al 24 for detailed overviews.
@@
  • lack of a thermodynamic limit due to the divergence of the electrostatic potential of the nuclei entering the 1e Hamiltonian $h$ used to define $U$.
The lack of 2ePC in 1eX2C has been shown to significantly impact the accuracy of molecular properties and spectra54, 23, thus calling for post-1eX2C treatment. The latter poses issues for extended systems, such as periodic systems. To avoid this problem, the application of 1eX2C to periodic systems55 must ad hoc replace the nuclear Coulomb potential by its Ewald counterpart. Proper application of X2C to extended systems also requires going beyond 1eX2C to define $U$ using 1-particle operators involving the Coulomb potential of electrons to ensure charge neutrality56.
The straightforward step past 1eX2C is to use the normal-ordered form of $H$, thereby defining the molecular mean-field X2C (mmfX2C)50, 57:
@@
  • Atomic-mean-field (AMF)-based methods, which correct for 2ePC using appropriately patched atomic densities and potentials obtained from 4C atomic (ensemble) Dirac-Hartree-Fock calculations. These methods include: the AMFI-X2C module66, the X2CAMF54, 67 scheme, and recent amfX2C and (extended) eamfX2C schemes23. Henceforth, we shall use the term AMF to refer to amfX2C, eamfX2C, and X2CAMF.
@@ 2 Formalism
Consider the RKB matrix form of the Dirac equation, also known as the modified Dirac equation, for a single particle in potential VV:70
$\begin{pmatrix}\mathbf{V}&\mathbf{T}\\ \mathbf{T}&\frac{\mathbf{W}}{4c^{2}}-\mathbf{T}\end{pmatrix}\begin{pmatrix}\mathbf{C}^{\text{L}}_{+}&\mathbf{C}^{\text{L}}_{-}\\ \mathbf{C}^{\text{S}}_{+}&\mathbf{C}^{\text{S}}_{-}\end{pmatrix}=\begin{pmatrix}\mathbf{S}&\mathbf{0}\\ \mathbf{0}&\frac{\mathbf{T}}{2c^{2}}\end{pmatrix}\begin{pmatrix}\mathbf{C}^{\text{L}}_{+}&\mathbf{C}^{\text{L}}_{-}\\ \mathbf{C}^{\text{S}}_{+}&\mathbf{C}^{\text{S}}_{-}\end{pmatrix}\begin{pmatrix}\bm{\epsilon}_{+}&\mathbf{0}\\ \mathbf{0}&\bm{\epsilon}_{-}\end{pmatrix}$, (5)
where $\bm{\epsilon}_{\pm}$ and $\mathbf{C}^{\text{L,S}}_{\pm}$ are the energies and corresponding AO coefficients, and $\mathbf{S}$ and $\mathbf{T}$ represent the identity and kinetic energy operators, respectively, in the L spinor basis. $\mathbf{V}$ and $\mathbf{W}$ are representations of the potential $V$ in the L and S spinor bases, respectively.
In the 1eX2C method, the transformation $U$ in Eq. 2 is defined to block diagonalize the core Dirac Hamiltonian
$\mathcal{H}=\begin{pmatrix}\mathbf{V}&\mathbf{T}\\ \mathbf{T}&\frac{\mathbf{W}}{4c^{2}}-\mathbf{T}\end{pmatrix}$ (6)
@@ 2.2 SAP-X2C
where $r_{A}=|\mathbf{r}-\mathbf{R}_{A}|$, and the effective nuclear charge function $Z^{\text{eff}}_{A}$ is defined on a fixed radial grid. Ref. 69 replaced the original prescription by a combination of the bare nuclear potential and the contribution due to the screening of the bare nucleus by other electrons:
@@ 3 Technical details
The 1eX2C and SAP-X2C Hamiltonians were implemented in the Massively Parallel Quantum Chemistry (MPQC) package77. The SAP basis set sap_grasp_large from Basis Set Exchange (BSE)78 is used for all SAP-X2C calculations in this work, as it is optimized with numerical Dirac-Coulomb Hartree Fock69. Note that the normalization convention for the SAP fitting basis is not the same as the convention for Gaussian AO basis sets, see Ref.69; the SAP definitions were incorporated into the Libint library’s public 2.10.0 release.74
The molecular geometries used for the comparison of HF energies in Figure 1 were optimized using Psi479 with PBE080 and def2-TZVP81 basis set. The molecular geometry of the Og2 molecule used for spinor energy assessment was taken from Ref.23. The dyall-ae3z82, 83, 84 basis set was used for all calculations in this work, except for the size-intensivity test, where dyall-2zp was used for efficiency. The physical constants defined in CODATA-202285 were used throughout, including the speed of light of 137.03599917697 $a_{0}E_{h}/\hbar$ obtained from the fine structure constant. The 4C-DCHF, amfX2C and eamfX2C calculations were performed using DIRAC12. The prerequisite atomic calculations needed for amf and eamfX2C were done using the average-of-configuration (AOC)86 mmfX2C with ground state electronic configurations taken from NIST Atomic Spectra Database87. The X2CAMF calculations were performed using socutils88. Since X2CAMF includes only 2eSO-PC and not 2eSR-PC, we are comparing against X2CAMF only for spin-orbit splitting energies, equilibrium bond distances, and harmonic frequencies.
@@ 4.1 Absolute Energies
Figure 1 shows errors in the total HF energies obtained with a variety of X2C approaches for a diverse set of small molecules containing light and heavy elements. Due to the lack of 2ePC, the 1eX2C errors clearly increase with $Z$. The amfX2C approach is highly accurate for most of the molecules presented, except for two outliers: RuC and PtC. eamfX2C, on the other hand, is highly accurate throughout; however, it is more expensive than amfX2C, as it involves the construction of a molecular 4C Fock matrix. SAP-X2C is a significant improvement over the 1eX2C and shows remarkably small errors even for heavy-element-containing systems.
@@
$\epsilon_{i}$ 4C-DCHF ($E_{h}$) amfX2C eamfX2C SAP-X2C 1eX2C sfX2C
@@ 4.2 Molecular Properties
Unlike the absolute energies and spin-orbit splittings, where SAP-X2C (while far superior to 1eX2C) never approached the performance of AMF-based X2C variants, SAP-X2C bond distances and harmonic vibrational frequencies were most accurate among all X2C variants for 3 and 2 molecules, respectively (out of 6). Surprisingly, amfX2C produced large errors for bond distances of Ag2 and Au2 ($10~\text{m}\AA$ ) and for the vibrational frequency of Au2 ($5~\text{cm}^{-1}$). While for most systems all X2C variants reproduce the reference bond distance to better than 1 milliangstrom, it is also intriguing to see the relatively wide distribution of the X2C equilibrium bond distances for At2 (Figure 2(f)). The observed excellent performance of SAP-X2C for equilibrium geometry and vibrational frequency computation makes it a prime candidate for inexpensive and accurate navigation of relativistic molecular PES.
Figure 2: Aligned potential energy surfaces of coinage and halogen dimers with 4C-DCHF and X2C-HF methods. $\Delta E\equiv E(R)-E(R_{\text{eq}})$.
@@ 4.3 Size-Intensivity of SAP-X2C Picture Change
The 1eX2C Hamiltonian (Eq. 2) does not have a thermodynamic limit due to the fact that the picture-change matrix $\mathcal{U}$ is obtained by diagonalization of the core Dirac Hamiltonian (Eq. 6). The latter contains the electrostatic potential of the nuclei alone, which diverges in the thermodynamic limit unless compensated by the electronic counterpart. This means that the 1eX2C picture-change matrix, and hence the 1eX2C Hamiltonian itself, do not have a thermodynamic limit. For periodic systems55 this can be mitigated by an ad hoc replacement of the nuclear electrostatic potential in the core Dirac Hamiltonian by the nuclear Ewald potential; however, for non-periodic systems or in the limit of an increasing unit cell, the fundamental issue remains.
@@ 5 Summary
We introduced a simple relativistic exact 2-component (X2C) Hamiltonian that accounts for two-electron picture-change effects using their free-atom model based on Lehtola’s Superposition of Atomic Potentials (SAP).68, 89 The SAP-X2C approach preserves the low-cost and technical simplicity of the popular 1-electron X2C (1eX2C) predecessor but is significantly more accurate and has a well-defined thermodynamic limit. Specifically:
@@
Given the promising performance, a far more extensive testing of SAP-X2C in its present form is clearly warranted. It may, however, be possible to improve the accuracy of SAP-X2C further by “tuning” the atomic model potentials for this purpose. This and other improvements will be explored in future work.
@@
The authors would like to acknowledge many insightful discussions with Trond Saue. This work was supported by the U.S. Department of Energy via award DE-SC0022327. The development of the Libint software library is supported by the Office of Advanced Cyberinfrastructure, National Science Foundation (Award OAC-2103738). The authors acknowledge Advanced Research Computing at Virginia Tech

### v2

@@ Abstract
We present a simple relativistic exact 2-component (X2C) Hamiltonian that models two-electron picture-change effects using Lehtola’s superposition of atomic potentials (SAP) [S. Lehtola, J. Chem. Theory Comput. 15, 1593–1604 (2019)]. The SAP-X2C approach keeps the low-cost and technical simplicity of the popular 1-electron X2C (1eX2C) predecessor, but is significantly more accurate and has a well-defined thermodynamic limit, making it applicable to extended systems (such as large molecules and periodic crystals). The assessment of the SAP-X2C-based Hartree-Fock total and spinor energies, spin-orbit splittings, equilibrium bond distances, and harmonic vibrational frequencies suggests that SAP-X2C is similar to the more complex atomic mean-field (AMF) X2C counterparts in its ability to approximate the 4-component Dirac-Hartree-Fock reference.
@@ 1 Introduction
Relativistic quantum chemistry1, 2, 3 is an approximation of quantum electrodynamics4 that combines Dirac 1-particle kinematics with the classical description of electromagnetic interactions in a many-particle system. By accounting for the dominant relativistic effects that are relevant for the chemistry and low-energy physics of atomistic matter and preserving the structure of nonrelativistic quantum many-body theory, accurate practical applications have become possible1, 5, 6, 7, 8, 9, 2, 10, 11, 12, 13, 14, 15. Despite the close resemblance to the nonrelativistic quantum chemistry, the appearance of positron-like degrees of freedom in the theory, such as bispinors (instead of spinors) and negative energy states, complicates both the formal and practical aspects of relativistic quantum chemistry. Thus, from the earliest days, there was keen interest in constructing electron-only (2-component, 2C) heuristics that accurately approximate the reference 4-component (4C) treatment. So-called quasi-relativistic 2C methods16, 17, 18, 10, 19, 20, 21, 22, 23, 24 include: the Foldy-Wouthuysen (FW)25 transformation, the Douglas-Kroll-Hess (DKH)26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36 methods, the zeroth-order regular approximation (ZORA)37, 38, 38, 39, Barysz-Sadlej-Snijders (BSS)40, 19, 41 method, and exact-two-component (X2C) methods42, 43 are generalized forms of the one-step normalized elimination of small component (NESC) approach prescribed by Dyall44, 45, 46, 47, 48 and self-consistent decoupling procedure49. Our work adds to the X2C family of methods42, 50, 51, 52, 43, 53, whose common trait is that the Fock-space Hamiltonian $H$ and other property operators expressed in a finite basis are rotated by a unitary matrix $U$ that is designed to block diagonalize the 1-particle Hamiltonian $h$ or its effective Fock-like equivalent (see Appendix A for notation). Subsequently, the rotated Hamiltonian is projected onto the upper (electron-like) components of the bispinor:
@@
The resulting X2C implementations differ in the details of these two design factors; see reviews by Liu18 and Wang et. al 24 for detailed overviews.
@@
  • non-size-intensive picture change and lack of a thermodynamic limit due to the divergence of the electrostatic potential of the nuclei entering the 1e Hamiltonian $h$ used to define $U$.
The lack of 2ePC in 1eX2C has been shown to significantly impact the accuracy of molecular properties and spectra54, 23, thus calling for post-1eX2C treatments. The latter poses issues for extended systems, such as periodic systems. To avoid this problem, the application of 1eX2C to periodic systems55 must ad hoc replace the nuclear Coulomb potential by its Ewald counterpart. Proper application of X2C to extended systems also requires going beyond 1eX2C to define $U$ using 1-particle operators involving the Coulomb potential of electrons to ensure charge neutrality56.
The straightforward step past 1eX2C is to use the normal-ordered form of $H$, thereby defining the molecular mean-field X2C (mmfX2C)51, 57:
@@
  • Atomic-mean-field (AMF)-based methods, which correct for 2ePC using appropriately patched atomic densities and potentials obtained from 4C atomic (ensemble) Dirac-Hartree-Fock calculations. These methods include: the AMFI-X2C module66, the X2CAMF54, 67 scheme, and the recent amfX2C and (extended) eamfX2C schemes23. Henceforth, we shall use the term AMF to refer to amfX2C, eamfX2C, and X2CAMF.
@@ 2 Formalism
Consider the RKB matrix form of the Dirac equation, also known as the modified Dirac equation,70 for a single particle in potential $V$ expressed in the orthonormal AO (OAO) basis:
$\begin{pmatrix}\mathbf{V}&\mathbf{T}\\ \mathbf{T}&\frac{\mathbf{W}}{4c^{2}}-\mathbf{T}\end{pmatrix}\begin{pmatrix}\mathbf{C}^{\text{L}}_{+}&\mathbf{C}^{\text{L}}_{-}\\ \mathbf{C}^{\text{S}}_{+}&\mathbf{C}^{\text{S}}_{-}\end{pmatrix}=\begin{pmatrix}\mathbf{1}&\mathbf{0}\\ \mathbf{0}&\frac{\mathbf{T}}{2c^{2}}\end{pmatrix}\begin{pmatrix}\mathbf{C}^{\text{L}}_{+}&\mathbf{C}^{\text{L}}_{-}\\ \mathbf{C}^{\text{S}}_{+}&\mathbf{C}^{\text{S}}_{-}\end{pmatrix}\begin{pmatrix}\bm{\epsilon}_{+}&\mathbf{0}\\ \mathbf{0}&\bm{\epsilon}_{-}\end{pmatrix}$, (5)
where $\bm{\epsilon}_{\pm}$ and $\mathbf{C}^{\text{L,S}}_{\pm}$ are the energies and corresponding OAO coefficients, and $\mathbf{1}$ and $\mathbf{T}$ represent the identity and kinetic energy operators, respectively, in the L orthonormal spinor basis. $\mathbf{V}$ and $\mathbf{W}$ are representations of the potential $V$ in the L and S spinor bases, respectively.
In the 1eX2C method, the transformation $U$ in Eq. 2 is defined to block diagonalize the core Dirac Hamiltonian.
$\mathcal{H}=\begin{pmatrix}\mathbf{V}&\mathbf{T}\\ \mathbf{T}&\frac{\mathbf{W}}{4c^{2}}-\mathbf{T}\end{pmatrix}$. (6)
@@ 2.2 SAP-X2C
where $r_{A}=|\mathbf{r}-\mathbf{R}_{A}|$, and the effective nuclear charge function $Z^{\text{eff}}_{A}$ is defined on a fixed radial grid by fitting to the reference potentials. Although the atomic potentials in SAP are local, they inherit the local description of exchange and correlation effects from the reference potential used for the fitting. Ref. 69 replaced the original prescription by a combination of the bare nuclear potential and the contribution due to the screening of the bare nucleus by other electrons:
@@
Here, we would like to point out that ZORA(MP, i.e., model-potential) by van Wüllen77 does share some connections with SAP-X2C; however,, the model-potential constructed in Ref.77 does not account 2ePC effects.
@@ 3 Technical details
The 1eX2C and SAP-X2C Hamiltonians were implemented in the Massively Parallel Quantum Chemistry (MPQC) package78. The SAP “basis” sap_grasp_large from Basis Set Exchange (BSE)79 is used for all SAP-X2C calculations in this work, as it is optimized with numerical Dirac-Coulomb Hartree Fock69. Note that the normalization convention for the SAP “basis” is not the same as the convention for Gaussian AO basis sets, see Ref.69; the SAP definitions were incorporated into the Libint library’s public 2.10.0 release.74
The molecular geometries used for the comparison of HF energies in Figure 1 were optimized using Psi480 with PBE081 and def2-TZVP82 basis set. The molecular geometry of the Og2 molecule used for spinor energy assessment was taken from Ref.23. The dyall-ae3z83, 84, 85 basis set was used for all calculations in this work, except for the size-intensivity test, where dyall-2zp was used for efficiency. The physical constants defined in CODATA-202286 were used throughout, including the speed of light of 137.03599917697 $a_{0}E_{h}/\hbar$ obtained from the fine structure constant. The 4C-DCHF, amfX2C and eamfX2C calculations were performed using DIRAC12. The prerequisite atomic calculations needed for amf and eamfX2C were done using the average-of-configuration (AOC)87 mmfX2C with ground state electronic configurations taken from NIST Atomic Spectra Database88. The X2CAMF calculations were performed using socutils89. Since X2CAMF includes only 2eSO-PC and not 2eSR-PC, we are comparing against X2CAMF only for spin-orbit splitting energies, equilibrium bond distances, and harmonic vibrational frequencies.
@@ 4.1 Absolute Energies
Figure 1 shows errors in the total HF energies obtained with a variety of X2C approaches for a diverse set of small molecules containing light and heavy elements. Due to the lack of 2ePC, the 1eX2C errors clearly increase with $Z$. The amfX2C approach is highly accurate for most of the molecules presented, except for two outliers: RuC and PtC. eamfX2C, on the other hand, is highly accurate throughout; however, it is more expensive than amfX2C, as it involves a singular construction of a molecular 4C Fock matrix. SAP-X2C is a significant improvement over the 1eX2C and shows remarkably small errors even for heavy-element-containing systems.
@@
$\epsilon_{i}$ 4C-DCHF ($E_{h}$) amfX2C eamfX2C SAP-X2C 1eX2C sf-1eX2C
@@ 4.2 Molecular Properties
Unlike the absolute energies and spin-orbit splittings, where SAP-X2C (while far superior to 1eX2C) never approached the performance of AMF-based X2C variants, SAP-X2C bond distances and harmonic vibrational frequencies were the most accurate among all X2C variants for 3 and 2 molecules, respectively (out of 6). Surprisingly, amfX2C produced large errors for bond distances of Ag2 and Au2 ($10~\text{m}\AA$ ) and for the vibrational frequency of Au2 ($5~\text{cm}^{-1}$). While for most systems all X2C variants reproduce the reference bond distances to better than 1 milliangstrom, it is also intriguing to see the relatively wide distribution of the X2C equilibrium bond distances for At2 (Figure 2(f)). The observed excellent performance of SAP-X2C for equilibrium geometry and vibrational frequency computation makes it a prime candidate for inexpensive and accurate navigation of relativistic molecular PES. Note that the differences in bond distances and vibrational frequencies resulting from different X2C Hamiltonians are relatively small compared to the residual errors of the electronic structure model itsel. This is illustrated in Table 6 that juxtaposes the equilibrium bond distances and harmonic vibrational frequencies obtained with with SAP-X2C and 1eX2C Hamiltonians using KS DFT (the hybrid PBE0 functional was used). The differences are relatively small relative to the HF-KS difference, but nevertheless becomes non-negligible for the heaviest systems.
Figure 2: Aligned potential energy surfaces (PESs) of coinage dimers (2(a),2(b),2(c)) and halogen dimers (2(d),2(e),2(f)) with 4C-DCHF and X2C-HF methods. $\Delta E\equiv E(R)-E(R_{\text{eq}})$.
@@
Equilibrium Bond Distance (Å) Harmonic Vibrational Frequency (cm-1)
Molecule 1eX2C SAP-X2C Experiment 1eX2C SAP-X2C Experiment
Cu2 2.2358 2.2356 2.21890 259.13 259.21 266.4990
Ag2 2.5685 2.5682 2.53091 181.97 181.94 192.492
Au2 2.4997 2.4994 2.47293, 91, 94, 95 186.50 186.24 190.993, 91, 94, 95
Br2 2.2805 2.28 2.28196, 97 332.59 333.04 325.3296, 97
I2 2.6742 2.6725 2.66598 215.69 216.59 214.5099, 100
At2 2.9873 2.9799 117.68 118.96
Table 6: Performance of 1eX2C and SAP-X2C Kohn-Sham with dyall-ae3z/PBE0 for coinage and halogen dimers compared to experimental values.
@@ 4.3 Size-Intensivity of SAP-X2C Picture Change
The 1eX2C Hamiltonian (Eq. 2) does not have a thermodynamic limit due to the fact that the picture-change matrix $\mathcal{U}$ is obtained by diagonalization of the core Dirac Hamiltonian (Eq. 6). The latter contains the electrostatic potential of the nuclei alone, which diverges in the thermodynamic limit unless compensated by the electronic counterpart. This means that the 1eX2C picture-change matrix, and hence the 1eX2C Hamiltonian itself, do not have a thermodynamic limit. For periodic systems55 this can be mitigated by an ad hoc replacement of the nuclear electrostatic potential in the core Dirac Hamiltonian by the nuclear Ewald potential. Such replacement is problematic on its own since the nuclear density is not charge neutral, thus the potential itself is not uniquely defined (although the asymptotic convergence with the supercell size can be fast; see Ref. 101). Furthermore, for non-periodic systems the fundamental issue remains.
@@ 5 Summary
We introduced a simple relativistic exact 2-component (X2C) Hamiltonian that accounts for two-electron picture-change effects using their free-atom model based on Lehtola’s Superposition of Atomic Potentials (SAP).68, 69 The SAP-X2C approach preserves the low-cost and technical simplicity of the popular 1-electron X2C (1eX2C) predecessor but is significantly more accurate and has a well-defined thermodynamic limit. Specifically:
@@
Given the promising performance, far more extensive testing of SAP-X2C in its current form is clearly warranted. However, it may be possible to further improve the accuracy of SAP-X2C by “tuning” atomic model potentials for this purpose. This and other improvements will be explored in future work.
{suppinfo}
Machine-readable XYZ files with Cartesian geometries for all species; description of the method used to construct crystal fragments used for size-intensivity tests in Section 4.3.
@@
The authors acknowledge many insightful discussions with Trond Saue. This work was supported by the U.S. Department of Energy via award DE-SC0022327. The development of the Libint software library is supported by the Office of Advanced Cyberinfrastructure, National Science Foundation (Award OAC-2103738). The authors acknowledge Advanced Research Computing at Virginia Tech